# Performance of Large Language Models in Numerical vs. Semantic Medical Knowledge: Benchmarking on Evidence-Based Q&As

### v2

@@ Abstract
Methods: To generate straightforward multi-choice questions and answers (QAs) based on evidence-based medicine (EBM), we used a comprehensive medical knowledge graph (encompassed data from more than 50,00 peer-reviewed articles) and created the “EBMQA”. EBMQA contains 105,000 QAs labeled with medical and non-medical topics and classified into numerical or semantic questions. We benchmarked this dataset on two state-of-the-art LLMs, Chat-GPT4 and Claude3-Opus. We evaluated the LLMs accuracy on semantic and numerical question types and according to sub-labeled topics. For validation, six medical experts were tested on 100 numerical EBMQA questions.
@@ 1 Introduction
To create a dataset that consists solely of EBM knowledge and is flexible enough to generate both semantic and numeric QA we used the Kahun knowledge graph- a clinically validated artificial intelligence tool that uses a medical, evidence-based knowledge graph. We have developed a methodology to generate QAs from this knowledge graph and created the EBMQA dataset. The dataset comprises 105,000 short multiple-choice questions, based on insights extracted from full-length articles, and aimed to test LLMs ability to assist physicians.
@@ 2.1.2 Questions Structure
All QAs were derived from Kahun’s knowledge graph. Each question was generated based on data from nodes and edges in the graph and consisted of three main entities: source (usually a disorder that is related to the target), target (usually a symptom or sign that is related to the source), and background (usually a relevant population related to the source). In this study, we refer to source, target, or background as entities. Additionally, the relation between entities (derived from data on the edges) determines the question type and the specific template that was used to generate the question and the answers.
@@ 2.2.1 QA selection and subanalysis
Due to the relatively different structure of semantic QA and its small number of QA, we analyzed all of them separately. Additionally, we subanalyzed this cohort by the number of different correct answers (one to five correct answers). Regarding numeric QA types and in the search for meaningful parameters that might influence LLM’s performance, the benchmark included QAs based on the following medical and non-medical sub-labels. Detailed descriptions of the QA selection criteria, sub-analyses, and the rationale behind each label used in this benchmark are provided in Appendix (Figure S1). Although different in the total number of QA per label, each label contains an identical number of selected QAs per sub-labeled entity. Additionally, each QA was selected once.
@@ 2.2.3 Evaluating LLM’s performance
  1. Accuracy - the total number of correct answers suggested by the LLM divided by the total answers suggested by the LLM, excluding IDK answers.
  2. Answer rate (AR) - the total number of both correct and wrong answers suggested by the LLM divided by the total answers suggested by the LLM, including IDK answers.
@@
2.2.4 Human and Prompt validation
To validate the benchmark, 100 numerical questions (questionnaire) were randomly selected. Two physicians and four medical students (in their clinical years) answered the questionnaire, first with the IDK possibility and then with guessing the answer to their IDK questions. Their accuracies with and without guessing were compared to LLMs. Additionally, to test both the effects of adding IDK as a possibility and changing the order of answers (including IDK), eight different prompts were tested: four prompts, each with a different order of possible answers, including IDK and four prompts each with different order of possible answers, without IDK.
@@ 2.2.5 Analysis and variables
All statistical analyses were performed using R-studio (version 4.2.2). Categorical variables were represented by a percentage while continuous variables were represented by mean and standard deviation (SD) if distrusted normally or by median and interquartile range (IQR) else. The cutoff for statistically significant results and confidence intervals (CI) were set to 95%. Proportions comparison was conducted using the Chi-Square test. For estimating the correlation of two quantitative variables, we employed the Spearman correlation.
@@ 3.2 Benchmark Analysis
3.2.1 Validation Analysis
Calude3 and GPT4 average accuracy rates, with the IDK option (64.25%,±3.95, 55.75%±1.71, respectively) or without it (59.25%±5, 53.25%±2.89, respectively), were higher than random (33%) or majority guessing (47%) though had lower average accuracy rates in comparing to humans with the IDK option (82.3%±2.82) or without it (78.2%±3.6) (Figure 2, Table S3).
@@ 3.2.2 Numeric QA subanalysis
Figure 3: Numeric QA accuracy and answer-rate sub-labels analysis: (A) Answer distribution, (B) Medical Discipline, (C) Medical Subject type, (D) QA type, (E) Disorders Prevalence, (F) Question length.
@@ 4 Discussion
The use of knowledge graphs for evaluating LLMs is gaining popularity [20, 21, 22]. Kahun’s structured knowledge graph enabled us to generate both semantic and numeric labeled QA, without utilizing advanced models [22]. Additionally, this unique knowledge graph allowed us to create a massive EBM dataset. Moreover, we embraced a data-driven approach in which distractors were based on sub-analysis distribution rather than specific or random values.
The EBMQA, which consists of 105,000 straightforward single-line QAs, was designed to mimic physicians’ strategy of breaking complex medical scenarios into less complicated problems, unlike medical licensing examination datasets which are typically complex-cases oriented [14, 23]. Additionally, the EBMQA addresses numeric/semantic data which is considered fundamental for physicians [24, 25] while dealing with data from articles and embracing the EBM approach [7], as opposed to the abstracted-based yes/no/maybe QAs in PubMedQA [15].
@@
While a recent benchmark analysis found that GPT4 outperformed Claude2 in nephrology questions [27], our inter-model examination reveals that Claude3 outperformed GPT4, including in medical disciplines such as neoplastic disorders (Figure 3B). Those results raise the need to constantly benchmark new LLMs as they continuously improve.
@@
Moreover, this comprehensive benchmark widens the medical scope and further supports both intra- and inter-model differences by exploring medical subjects: as Claude3 favors “Lab test” (66%) and struggles with “Disorders” (60%), GPT4 prefers “Imaging and Procedures” (61%) rather than “Lab test” (53%).
@@
In terms of prompt engineering, our validation test showed relatively small SDs in prompt accuracy which supported our prompt stability. Moreover, although insignificant, the IDK prompt yielded higher accuracy and, therefore was used.

### v3

@@ Abstract
Methods: To generate straightforward multi-choice questions and answers (QAs) based on evidence-based medicine (EBM), we used a comprehensive medical knowledge graph (encompassed data from more than 50,00 peer-reviewed articles) and created the “EBMQA”. EBMQA contains 105,222 QAs labeled with medical and non-medical topics and classified into numerical or semantic questions. We benchmarked this dataset on two state-of-the-art LLMs, Chat-GPT4 and Claude3-Opus. We evaluated the LLMs accuracy on semantic and numerical question types and according to sub-labeled topics. For validation, six medical experts were tested on 100 numerical EBMQA questions.
@@ 1 Introduction
To create a dataset that consists solely of EBM knowledge and is flexible enough to generate both semantic and numeric QA we used the Kahun knowledge graph- a clinically validated artificial intelligence tool that uses a medical, evidence-based knowledge graph. We have developed a methodology to generate QAs from this knowledge graph and created the EBMQA dataset. The dataset comprises 105,222 short multiple-choice questions, based on insights extracted from full-length articles, and aimed to test LLMs ability to assist physicians.
@@ 2.1.2 Questions Structure
All QAs were derived from Kahun’s knowledge graph. Each question was generated based on data from nodes and edges in the graph and consisted of three main entities: source (usually a disorder that is related to the target), target (usually a symptom or sign that is related to the source), and background (usually a relevant population related to the source). In this study, we refer to source, target, or background as entities. Additionally, the relation between entities (derived from data on the edges) determines the question type and the specific template that was used to generate the question and the answers.Further explanation regarding template creation provided in the Appendix.
@@ 2.2.1 QA selection and subanalysis
Due to the relatively different structure of semantic QA and its small number of QA, we analyzed all of them separately. Regarding numeric QA types and in the search for meaningful parameters that might influence LLM’s performance, the benchmark included QAs based on three medical (Medical subject type, Medical Discipline and Prevalence) and three non-medical sub-labels (QA types, Question length and Answers distribution) as further detailed in Appendix (Figure S1). All QAs were selected randomly, and although the total number of QAs per label varied, each label contains an identical number of selected QAs per sub-labeled entity. Additionally, each QA was selected only once.
@@ 2.2.3 Evaluating LLM’s performance
  1. Accuracy - for both semantic and numeric QAs, the total number of correct answers suggested by the LLM divided by the total answers suggested by the LLM, excluding IDK answers.
  2. Answer rate (AR) - for both semantic and numeric QAs, the total number of both correct and wrong answers suggested by the LLM divided by the total answers suggested by the LLM, including IDK answers.
@@
2.2.4 Prompt sensitivity analysis
To test both the effects of adding IDK as a possible answer and changing the order of answers (including IDK), eight different prompts were tested on 100 numerical questions that were randomly selected (questionnaire): four prompts, each with a different order of possible answers, including IDK, and four prompts, each with different order of possible answers, without IDK.
2.2.5 Human validation
To validate the QAs in the EBMQA, two physicians and four medical students (in their clinical years) answered the questionnaire, first with the IDK possibility and then with guessing the answer to their IDK questions. Their accuracies with and without guessing were compared to LLMs’.
@@ 2.2.6 Analysis and variables
All statistical analyses were performed using R-studio (R version 4.2.2). Categorical variables were represented by a percentage while continuous variables were represented by mean and standard deviation (SD) if distributed normally or by median and interquartile range (IQR) else. The cutoff for statistically significant results was set at was set at alpha= 0.05 and confidence intervals (CI) were set to 95%. Proportions comparison was conducted using the ”Proportion test”. For estimating the correlation of two quantitative variables, we employed the Spearman correlation.
@@ 3.2 Benchmark Analysis
3.2.1 Prompt sensitivity analysis
[…]
3.2.2 Human Validation
Calude3 and GPT4 average accuracy rates, with the IDK option or without it were higher than random (33%) or majority guessing (47%) though had lower average accuracy rates in comparing to humans with the IDK option ($82.3\%\pm 2.82$) or without it ($78.2\%\pm 3.6$) (Figure 2, Table S3).
@@ 3.2.3 Numeric QA subanalysis
Figure 3: Numeric QA accuracy and answer-rate sub-labels analysis: (A) Answer distribution, (B) Medical Discipline, (C) Medical Subject type, (D) QA type, (E) Disorders Prevalence, (F) Question length.Red asterisks represent proportion p-values: .05$<$ *$<$.01, ***$<$.0001
@@ 4 Discussion
The use of knowledge graphs for evaluating LLMs is gaining popularity [20, 21, 22]. Kahun’s structured knowledge graph enabled us to generate both semantic and numeric labeled QA, without utilizing advanced models [22]. Our QAs generation process, which relies on templates designed to fit a source-target-background graph structure can be applicable to other graphs with a similar structure. Additionally, this relatively large knowledge graph allowed us to create a massive EBM dataset. Moreover, we embraced a data-driven approach in which distractors were based on sub-analysis distribution rather than specific or random values
The EBMQA, which consists of 105,222 straightforward single-line QAs, was designed to mimic physicians’ strategy of breaking complex medical scenarios into less complicated problems, unlike medical licensing examination datasets which are typically complex-cases oriented [14, 23]. Additionally, the EBMQA addresses numeric/semantic data which is considered fundamental for physicians [24, 25] while dealing with data from articles and embracing the EBM approach [7], as opposed to the abstracted-based yes/no/maybe QAs in PubMedQA [15].
@@
A recent benchmark analysis, focused on nephrology QAs only, found that GPT4 outperformed Claude2 [27]. Although our inter-model examination did not include a direct nephrology compression, due to a different classification method, it reveals that generally Claude3 outperformed GPT4 and specifically in a variety of medical disciplines such as neoplastic disorders, nervous system and more (Figure 3B). Those results raise the need to constantly benchmark new LLMs as they continuously improve.
@@
[…]
@@
In terms of prompt engineering, our sensitivity analysis showed relatively small SDs in prompt accuracy which supported our prompt stability. Additionally, although insignificant, the IDK prompt yielded higher accuracy and, therefore was used. Moreover, changing the order of the distractors did not significantly affect the performances of the LLMs.